### v1

# Confidence-weighted integration of human and machine judgments for superior decision-making

Abstract
Large language models (LLMs) have emerged as powerful tools in various domains. Recent studies have shown that LLMs can surpass humans in certain tasks, such as predicting the outcomes of neuroscience studies [1]. What role does this leave for humans in the overall decision process? One possibility is that humans, despite performing worse than LLMs, can still add value when teamed with them. A human and machine team can surpass each individual teammate when team members’ confidence is well-calibrated and team members diverge in which tasks they find difficult (i.e., calibration and diversity are needed). We simplified and extended a Bayesian approach to combining judgments using a logistic regression framework that integrates confidence-weighted judgments for any number of team members. Using this straightforward method, we demonstrated in a neuroscience forecasting task [1] that, even when humans were inferior to LLMs, their combination with one or more LLMs consistently improved team performance. Our hope is that this simple and effective strategy for integrating the judgments of humans and machines will lead to productive collaborations.
1 Introduction
Modern environments increasingly stretch our ability to process the vast amounts of information available to us [2, 3]. In contrast, machine systems can often take advantage of vast information resources [4, 5, 6, 7]. As machines reach superhuman performance levels [6, 8, 9], one concern is whether machines will supplant human judgment in critical areas [10, 11].
One potential solution is forming human-machine teams in which judgments from humans and machines are integrated [12, 13]. It might be possible that humans can contribute to and make the overall team better even when their performance is worse on average than their machine teammates.
We will evaluate this possibility in a knowledge-intensive task where large language models (LLMs) surpass humans in predicting the outcomes of neuroscience studies [1]. Human-LLM teaming combines the individual judgments of humans and machines. Complementarity is realized when a team’s performance improves beyond that of either teammate alone [12, 13]. We investigate whether human-LLM teams outperform LLMs even when humans have inferior performance compared to LLMs. There are two key conditions for team complementarity to be fulfilled [1, 12, 14]. The first requirement is calibration of confidence. This implies that when humans and LLMs have a higher degree of confidence in their judgments, the accuracy of those judgments tends to be greater [1]. The second requirement is classification diversity among team members. Diversity holds when the errors in classification made by humans and LLMs are not the same [1].
Previous work [12] has explored the conditions for complementarity in the context of object recognition. Humans outperformed machines in the classification of natural images, raising the question of whether a combined approach could achieve superhuman performance. They developed a Bayesian model that integrates the judgments of humans and machines. With this approach, human-machine complementarity was observed. However, the combination model is computationally expensive and challenging to extend to additional teammates. Ideally, a model that combines the judgments of humans and machines should be adaptable and scalable, easily interpretable, and allow for any number of teammates.
Here, we aim to offer this ideal solution to human-machine teaming while evaluating complementarity in a knowledge-intensive task that is not based on perceptual judgment and, critically, where humans were surpassed by machine systems. Foreshadowing our results, we find support for effective human-LLM teaming through our novel and resource-efficient procedure. Our procedure comprises a logistic-regression-based strategy that provides confidence-weighted integration of teammates’ predictions for any number of team members. Our approach is particularly well-suited for combining human and machine judgments and assessing their contribution in predictive tasks.
2 Methods
Figure 1: Assessing Humans and LLMs using BrainBench [1]. a) The benchmark consisted of test cases constructed from the Journal of Neuroscience abstracts. Abstracts consist of background, methods, and results. The test-taker chose which of two versions of the abstract was the original version. The altered version maintained coherency while significantly altering the results. The 100 test cases considered here were constructed by GPT-4 with human oversight and quality control. b) An example test case. Humans were instructed to select which version of the abstract was the original by clicking on either blue or green text to select that set of options. Test cases varied in the numbers of alternatives, but a single click will choose all options of the same color. After their choice, humans indicated their confidence. LLMs chose the version of the abstract that had the lower perplexity and their confidence was assessed by the absolute difference in perplexity of the two options.
Our contribution relies on previous efforts that developed BrainBench [1] to assess the capacity of humans and LLMs to predict the outcomes of neuroscience studies. The benchmark includes test cases based on abstracts from the Journal of Neuroscience. Each test case contains an original abstract and an altered version (Figure 1), LLMs outperform humans by a large margin in test cases created either by expert neuroscientists or by prompting GPT-4 to create test cases [1]. Since substantial differences in performance may preclude complementarity [12], we used the GPT-4 generated test cases because the performance difference between humans and LLMs was smaller for these test cases, though LLMs were still clearly superior to humans on these test items.
2.1 Test cases
BrainBench [1] is a benchmark that includes 100 test cases generated by GPT-4 (Azure OpenAI API; version 2023-05-15). These test cases were created from abstracts in the Journal of Neuroscience published in 2023. These abstracts are categorized into five sections: Behavioral/Cognitive, Systems/Circuits, Development/Plasticity/Repair, Neurobiology of Disease, and Cellular/Molecular. Each test case contains a published abstract and an altered version produced by GPT-4 (see details in [1, Dataset Creation]). These modifications, though minimal, significantly change the results—for instance, by changing the roles of brain regions or reversing a result’s direction (e.g., from “decreases” to “increases”). The altered abstracts remain logically coherent despite the changes. The BrainBench task is to identify the correct study outcome by choosing between the original abstract and its altered counterpart.
2.2 Procedure for human participants
One hundred seventy-one neuroscience experts were recruited to complete an online study (see details in [1, Evaluations]). Each participant evaluated three out of the 100 test cases. Two versions of an abstract were presented, one with the actual results and one that was altered (Figure 1). Participants chose the version they believed to be the original and rated their confidence using a slider bar. After applying several exclusion criteria, the 171 participants yielded 503 observations (2 to 9 instances per test case).
2.3 Procedure for LLMs
We considered Llama 2 chat models with 7B, 13B, and 70B weights [15]. LLMs chose the version of the abstract with the lower perplexity (PPL). Confidence was calculated as the absolute difference in PPL between the original and altered versions of the abstract (Figure 1).
2.4 Bayesian combination model
Human and LLM judgments were combined by adapting a Bayesian framework for human-machine complementarity [12]. The problem setting for combining two team members, human and LLM, is as follows: Let $N$ denote the number of test cases to be analyzed with two possible choices, “0: first option” or “1: second option”. The ground truth labels of the original abstracts are $z\in\{0,1\}^{N}$. For the human classifier, the predicted labels $y\in\{0,1\}^{N}$ and their corresponding confidence ratings $r\in\{0,\ldots,R-1\}^{N}$ with “0: lowest possible confidence” and “$R-1$: highest possible confidence” are given. For the LLM classifier, the PPL $q\in\mathbb{R}_{+}^{N\times 2}$ reflects a measure of uncertainty [1]. We used probability confidence scores, $\pi=\text{Softmax}(-q)$. The first step of this model is to generate correlated probability confidence scores for human and machine classifiers using a bivariate normal distribution,
$\left(\begin{matrix}\pi_{H}\\ \pi_{M}\end{matrix}\right)\sim\mathcal{N}\left(\left(\begin{matrix}\mu_{H}\\ \mu_{M}\end{matrix}\right),\left(\begin{matrix}\sigma_{H}^{2}&\sigma_{H}\sigma_{M}\rho_{HM}\\ \sigma_{H}\sigma_{M}\rho_{HM}&\sigma_{M}^{2}\end{matrix}\right)\right)\,.$ (1)
The means of the underlying distribution, $\mu_{H}$ and $\mu_{M}$, depend whether the label of test case $i$, $z_{i}$, is correct or not, i.e.,
$\mu_{H,i,j}$ $=b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,,$
$\mu_{M,i,j}$ $=b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,,$
with $\mathcal{Z}_{i}=\{x\,|\,x=z_{i}\}$. Note that the scalar parameters $a_{H}$, $a_{M}$, $b_{H}$, $b_{M}$, $\sigma_{H}$, $\sigma_{M}$, and $\rho_{HM}$ in Eq. (1) are learned from data. The parameter $\rho_{HM}$ from the covariance matrix learns the correlation between the human and machine classifiers. In the case of the machine classifier, $\pi_{M}$ is compared to the empirical probability confidence scores, $\pi$. Then, for the human classifier, $\pi_{H}$ is a latent variable that is used to calculate classifications
$y_{H}\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)\,,$
where $\tau$ denotes a temperature parameter, usually small, that helps convergence [12]. The predicted classifications, $y_{H}$, are compared to actual human predictions, $y$. The confidence ratings are casted via an ordered logistic model [16] that maps the continuous probability confidence scores, $\pi_{H}$, to an ordinal confidence rating, $r_{H}$. This means,
$r_{H}\sim\text{OrderedLogistic}(\pi_{H},c,\delta)\,,$
where the parameters $c\in\mathbb{R}_{+}^{R-1}$ are the breakpoints of the intervals that map $\pi_{H}$ into $r_{H}$, and $\delta$ is a scalar that controls the sharpness of the rating probability curves. Finally, $r_{H}$ is compared to the empirical human confidence ratings, $r$. See supplementary information for implementation details.
2.5 Logistic combination model
We introduce a logistic regression approach that combines the judgments of any number of teammates. The logistic combination model follows the principles of the Bayesian combination model, but is formulated within an easier to implement and extend regression framework. In its most basic form, which we consider here, each teammate is captured by a single predictor in the regression model. The value of the predictor on a trial depends on the teammate’s choice and their confidence. In particular, the magnitude of the predictor is the teammate’s confidence on that trial (i.e., confidence-weighted integration) and the sign is determined by the teammate’s choice. In general, the fitted beta weight for a teammate will reflect their accuracy and calibration.
As in the Bayesian combination model, $y$, $r$, $q$ are given. For the human classifier, the confidence ratings were of the form: $r\in\{1,\ldots,100\}^{N}$. In the case of LLMs, the absolute difference in PPL $|\Delta q_{i}|\in\mathbb{R}_{+}^{N}$ corresponds to model confidence [1].
The logistic function for the $i$-th test case is of the form:
$p_{x}=\frac{1}{1+e^{-\beta^{\top}x}}\,,$
where $p_{x}$ is the predicted probability of the arbitrarily assigned first option, and the evidence is
$\beta^{\top}x=\beta_{I}+\beta_{H}\cdot x_{H}+\beta_{M}\cdot x_{M}\,.$ (2)
The fitted weights $\beta_{I}$, $\beta_{H}$, and $\beta_{M}$ correspond to the intercept, and human and machine teammates, respectively. The term $x_{k}$ is the signed confidence for team member $k$. For human participants, $x_{H}$ is $r_{i}$ if $y_{i}=0$ and $-r_{i}$ otherwise. Thus,
$x_{H}=r_{i}(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,,$
where $\mathbf{1}_{\mathcal{A}}(y_{i})$ is an indicator function with $\mathcal{A}=\{x\,|\,x=1\}$. Hence, $\mathbf{1}_{\mathcal{A}}(x)=1$ if $x\in\mathcal{A}$, and $\mathbf{1}_{\mathcal{A}}(x)=0$ otherwise. For LLM team members, it reads
$x_{M}=|\Delta q_{i}|(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,.$
This model can be easily expanded by including additional terms to Eq. (2). For example, a third fitted weight could be included for an interaction term $x_{H}x_{M}$. Likewise, polynomial regression could be used to include $x_{H}^{2}$ and $x_{M}^{2}$, and corresponding fitted weights. This approach can be readily extended to more than 2 team members by using multinomial regression.
2.6 Cross-validation procedure
Given the 503 observations for the 100 test cases, we performed a leave-one-out cross-validation (LOOCV) [17, 18] that provides the best bias-variance trade-off for small datasets. Consider the evaluation of the $i$-th test case. With this procedure, we removed all instances of test case $i$ (between 2 and 9) leaving the remaining 99 test cases with all their instances (between 494 and 501) to train the classifier teams. For testing, we utilized all the instances of test case $i$. This was repeated for all 100 test cases, yielding 503 predictions. Note that for individual teammates the evaluation comprised only the testing phase. When the predicted labels of any team or individual teammate in this study were randomly shuffled, the LOOCV accuracy dropped approximately to chance level (i.e., 50%).
3 Results
Figure 2: Conditions for effective collaboration between human experts and LLMs were satisfied. a) When LLMs were confident in their BrainBench judgments, they were more likely to be correct. Confidence ratings were sorted into equal bins, and the mean accuracy for each bin was plotted. The positive slope of the black regression lines for Llama2 chat models (7B, 13B, and 70B) indicates well-calibrated confidence, meaning higher confidence correlated with higher accuracy. b) LLMs surpassed human experts on BrainBench overall. Error bars represent standard error of the mean using a binomial model. c) Item difficulty Spearman correlations among LLMs and human experts. For LLMs, difference in perplexity between incorrect and correct abstracts was used to determine the relative difficulty of test cases. Mean accuracy was used for human experts. LLMs align more with each other than humans, which implies human-machine teams will be diverse.
Figure 3: Performance of all possible teams using the confidence-weighted logistic combination model. Adding a human to a team with one or more machines (blue points) always has a benefit (green points). Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 evaluations. Error bars represent standard error of the mean using a binomial model.
We first assessed the conditions for effective collaboration (i.e., complementarity), namely calibration of confidence and classification diversity among team members. Both humans and LLMs were calibrated in that accuracy was positively correlated with confidence (Figure 2a). Diversity held in that LLMs and humans differed on which test items led to errors (Figure 2c).
In terms of accuracy (Figure 2b), LLMs numerically surpassed humans by a small margin ($t(2)=5.20$, $P<0.05$). Thus, we can consider whether humans can benefit teams consisting of machines that perform comparably or better.
Of primary interest was whether teams including humans performed better than LLM only teams. All 15 possible team combinations, ranging from individual teammates to a 4-way human-LLM team, were considered (Figure 3). Adding a human teammate to LLM-only teams always improved the team’s performance ($t(6)=8.22$, $P<0.0001$).
Our confidence-weighted logistic regression approach follows from the principles of a Bayesian combination model that fosters human-machine complementarity [12]. One questions is how well our logistic regresion approach compares to the Bayesian approach. Our confidence-weighted regression model compared favorably (Figure 4) to the Bayesian model when evaluated on the three human-LLM teams for which the Bayesian model is intended to apply ($t(2)=4.08$, $P<0.01$). This success is impressive given that the regression approach takes seconds to compute on a current desktop whereas the Bayesian approach would take 12-13 days.
Figure 4: Comparison between Bayesian and confidence-weighted logistic combination models for human-LLM teams. Each data point corresponds to the average across 503 evaluations. Llama2 chat 7B, 13B, and 70B models were considered. Error bars represent standard error of the mean using a binomial model. The confidence-weighted logistic combination model more effectively integrated human and machine judgments.
4 Discussion
Can humans team effectively with LLMs when the humans perform worse? We developed a confidence-weighted regression approach that can integrate judgments from any number of teammates. Using this method and testing on a forecasting benchmark (Figure 1), we found that human-LLM teams achieve complementarity, that is their combined performance bests that of either teammate alone (Figure 3). Complementarity was achieved because two critical conditions were satisfied, namely confidence was well-calibrated and classification diversity held among teammates (Figure 2). Strikingly, every combination of LLMs benefited from adding a human to the team (Figure 3).
Our approach was informed by a Bayesian method for combining judgments of humans and machines [12]. Our approach has a number of advantages, including ease of implementation, very fast runtime, an interpretable solution, and readily extendable to any number of teammates. Surprisingly, our confidence-weighted regression approach performed better than the Bayesian approach (Figure 4). One possibility is that the discretization of continuous confidence measures, which the Bayesian model requires, limited its performance. Perhaps an alternative formulation would perform better. Unfortunately, reformulating the Bayesian model and properly implementing it requires substantial effort and expertise. In contrast, because our confidence-weighted integration model is formulated within a regression framework, it is straightforward to extend to include other factors. For example, nonlinear relationships (e.g., polynomial terms) between confidence-weighted predicitons and outcomes could be considered.
While we selected three LLMs with superhuman performance on BrainBench, these LLMs are not the highest performing models on this benchmark [1]. Our choice was deliberate because a vastly superior teammate may hinder complementarity. In the limit, a teammate who is never wrong does not need to be part of a team. This limiting condition may become more prevalent should LLMs continue to improve and, therefore, diminish the benefits of human-LLM teaming. For the foreseeable future, we suspect there will be tasks for which humans and LLMs can effectively team. Moreover, our method for integrating the judgments of teammates is not limited to human-LLM teams. Instead, the method is general and applies to any set of agents (natural or artificial) that can report how confident they are in their decisions.
This study explored the possibility of a collaborative approach between humans and LLMs for superior decision-making in predicting neuroscience outcomes. Our confidence-weighted regression method effectively combined human and LLM judgments because teammates fulfilled the conditions of well-calibrated confidence and classification diversity. Our results suggest there is a place for humans in teams with machines even when the machines perform better. We hope our work facilitates successful collaborations between humans and machines in addressing important challenges.
5 Data availability
The human participant data and LLM perplexity scores utilized in this study have been previously reported in [1], and are publicly available at
https://github.com/braingpt-lovelab/BrainBench .
6 Code availability
All computer code associated with this work including combination model implementations, team evaluations, and analyses are publicly available at
https://github.com/braingpt-lovelab/haico .
Acknowledgments
This work was supported the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), and a Royal Society Wolfson Fellowship (18302) to B.C.L.
Supplementary information
Implementation of Bayesian combination model
Algorithm 1 Bayesian Combination Model
1:Set priors: $\alpha=1$, $a_{M}\sim\mathcal{N}(0,10)$, $b_{M}\sim\mathcal{N}(0,10)$, $\sigma_{M}\sim\text{Uniform}(0,15)$, $a_{H}\sim\mathcal{N}(0,10)$, $b_{H}=0$, $\sigma_{H}=1$, $\rho\sim\text{Uniform}(-1,1)$, $\tau=0.05$, $c\sim\text{Uniform}(0,1)$ with $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2$, and $\delta\sim\text{Uniform}(0,100)$.
2:start training:
3:     $p\sim\text{Dirichlet}(\alpha)$
4:     $z\leftarrow\text{argmax}(p)$$\triangleright$ Compare to actual data
5:     $\mu_{M,i,j}\leftarrow b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
6:     $\mu_{H,i,j}\leftarrow b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
7:     $\pi_{M}\sim\mathcal{N}\left(\mu_{M},\,\sigma_{M}\right)$triangleright Compare to actual data
8:     $\pi_{H}\sim\mathcal{N}\left(\mu_{H}+\rho\,\sigma_{H}\left(\frac{\pi_{M}-\mu_{M}}{\sigma_{M}}\right),\,\sqrt{1-\rho^{2}}\,\sigma_{H}\right)$
9:     $y\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)$triangleright Compare to actual data
10:     $r\sim\text{OrderedLogistic}(\pi_{H},c,\delta)$triangleright Compare to actual data
11:end training
12:start testing:
13:     $p\sim\text{Dirichlet}(\alpha)$
14:     $z\leftarrow\text{argmax}(p)$
15:     $\mu_{M,i,j}\leftarrow b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
16:     $\mu_{H,i,j}\leftarrow b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
17:     $\pi_{M}\sim\mathcal{N}\left(\mu_{M},\,\sigma_{M}\right)$triangleright Compare to actual data
18:     $\pi_{H}\sim\mathcal{N}\left(\mu_{H}+\rho\,\sigma_{H}\left(\frac{\pi_{M}-\mu_{M}}{\sigma_{M}}\right),\,\sqrt{1-\rho^{2}}\,\sigma_{H}\right)$
19:     $y\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)$triangleright Compare to actual data
20:     $r\sim\text{OrderedLogistic}(\pi_{H},c,\delta)$triangleright Compare to actual data
21:end testing
Algorithm 1 illustrates the inference procedure of the Bayesian combination model, which comprises two stages: estimation of parameters (training) and prediction of label probabilities (testing). For the estimation of parameters, the given data is: true labels $z$, probability confidence scores $\pi_{M}$, and human classification $y$ and confidence ratings $r$. We assumed that all human participants shared the same set of parameters ($a_{H}$ , $b_{H}$ , $\sigma_{H}$, $c$, $\delta$, and $\tau$). Human confidence ratings on the slider bar were mapped to range between 1 and 100. A wide range is computationally expensive, thus, we aggregate it into three levels: “0: low confidence”, “1: moderate confidence”, and “2: high confidence”. Then, the aggregated confidence rating used for analysis, $r\in\{0,1,2\}^{N}$, reads
[…]
The utilized cutpoints (i.e., 33 and 66) produced a good agreement between confidence and accuracy. Among the evaluations of human participants, “low” had $63.2\%$ average accuracy ($n=174$), “moderate” had $66.5\%$ ($n=185$), and “high” had $81.9\%$ ($n=144$). To estimate the posterior over the underlying parameters, a No-U-Turn Sampler (NUTS) for Markov chain Monte Carlo (MCMC) [19] was used with $n_{w}=500$ warmup steps, $n_{c}=3$ chains, and $n_{s}=25$ samples. For the prediction of label probabilities the given data is the same as before but $z$, that is now latent. The prior distributions were set according to [12]. The correlation between team members was $\rho\sim\mathrm{Uniform}(-1,1)$. For the LLM classifier, $a\sim\mathcal{N}(0,10)$, $b\sim\mathcal{N}(0,10)$, and $\sigma\sim\mathrm{Uniform}(0,15)$. For human classifications, $a$ was the same but $b$ and $\sigma$ were adjusted to $b=0$ and $\sigma=1$ for the purpose of identifiability [12]. In addition, we used $\delta\sim\mathrm{Uniform}(0,100)$ for the scaling parameter and uniform priors on the cutpoints, $c\sim\mathrm{Uniform}(0,1)$, with the constraint that the cutpoints are ordered (i.e. $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2)$. Finally, we set $\tau=0.05$ for best convergence results [12].