### v2

# Confidence-weighted integration of human and machine judgments for superior decision-making

Abstract
Large language models (LLMs) have emerged as powerful tools in various domains. Recent studies have shown that LLMs can surpass humans in certain tasks, such as predicting the outcomes of neuroscience studies. What role does this leave for humans in the overall decision process? One possibility is that humans, despite performing worse than LLMs, can still add value when teamed with them. A human and machine team can surpass each individual teammate when team members’ confidence is well-calibrated and team members diverge in which tasks they find difficult (i.e., calibration and diversity are needed). We simplified and extended a Bayesian approach to combining judgments using a logistic regression framework that integrates confidence-weighted judgments for any number of team members. Using this straightforward method, we demonstrated its effectiveness in both image classification and neuroscience forecasting tasks. Even when humans performed worse than machines, combining human judgments with one or more machines consistently improved overall team performance. Our hope is that this simple and effective strategy for integrating the judgments of humans and machines will lead to productive collaborations.
Keywords
Human-AI Collaboration, Large Language Models (LLMs), Neuroscience, Decision Making
1 Introduction
Modern environments increasingly stretch our ability to process the vast amounts of information available to us [1, 2]. In contrast, machine systems can often take advantage of vast information resources [3, 4, 5, 6]. As machines reach superhuman performance levels [5, 7, 8], one concern is whether machines will supplant human judgment in critical areas [9, 10].
One potential solution is forming human-machine teams in which judgments from humans and machines are integrated [11, 12, 13]. It might be possible that humans can contribute to and make the overall team better even when their performance is worse on average than their machine teammates.
We will begin evaluating this possibility in an object recognition task where human and machine performance overlap according to experimental conditions [13]. Human-machine teaming combines the individual judgments of humans and machines. We will then evaluate a knowledge-intensive task in which large language models (LLMs) surpass humans in predicting the outcomes of neuroscience studies [14], posing a real challenge for effective team collaboration. Complementarity is realized when a team’s performance improves beyond that of either teammate alone [13, 12]. We investigate whether human-LLM teams outperform LLMs even when humans have inferior performance compared to LLMs. There are two key conditions for team complementarity to be fulfilled [14, 13, 15]. The first requirement is calibration of confidence. This implies that when humans and LLMs have a higher degree of confidence in their judgments, the accuracy of those judgments tends to be greater [14]. The second requirement is classification diversity among team members. Diversity holds when the errors in classification made by humans and LLMs are not the same [14].
Previous work [13] has explored the conditions for complementarity in the context of object recognition. Humans outperformed machines in the classification of natural images with low levels of noise, raising the question of whether a combined approach could achieve superhuman performance. They developed a Bayesian model that integrates the judgments of humans and machines. With this approach, human-machine complementarity was observed. However, the combination model is computationally expensive and challenging to extend to additional teammates. Ideally, a model that combines the judgments of humans and machines should be adaptable and scalable, easily interpretable, and allow for any number of teammates.
Here, we aim to offer this ideal solution to human-machine teaming while evaluating complementarity in an object recognition task, and a knowledge-intensive task that is not based on perceptual judgment. Critically, in both scenarios, humans were surpassed by machine systems. Foreshadowing our results, we find support for effective human-machine teaming through our novel and resource-efficient procedure. Our procedure comprises a logistic-regression-based strategy that provides confidence-weighted integration of teammates’ predictions for any number of team members. Our approach is particularly well-suited for combining human and machine judgments and assessing their contribution in predictive tasks.
2 Results
In this study, we explore whether humans can contribute to decisions when machine models, such as LLMs, are superior to them. We developed a logistic-regression-based method that integrates a weighted average of judgments from teammates, whether humans or machines. The proposed approach adheres to similar previously reported principles [13], offering a number of advantages: it is easy to use, flexible, and resource-efficient (details can be found in the Methods section). We evaluate our method in an object recognition task where effective human-machine collaboration has been demonstrated [13]. We then shift our focus to BrainBench [14] because LLMs significantly outperform human experts, posing a real challenge for team collaboration.
Figure 1: Performance of the confidence-weighted logistic combination model in a noisy object recognition task [13]. Accuracy results on high levels of image noise ($\Omega=125$) with the logistic combination model. Adding a human to a team with one or more machines (blue points) always has a benefit (green points). Each data point corresponds to the average across 7 239 image evaluations. Error bars represent standard error of the mean using a binomial model.
Performance is improved when a human collaborates in a machine-only team
We first assessed the performance of human-machine teams in the classification of noisy natural images [13]. The images were distorted by phase noise at each spatial frequency, where the phase noise was uniformly distributed in the interval $[-\Omega,\Omega]$ [16]. We considered two noise levels: images distorted by low ($\Omega=80$) and high ($\Omega=125$) noise. In the case of low noise, machines are surpassed by humans ($t(4)=-4.77$, $P<0.01$). The Bayesian combination model demonstrated human-machine complementarity (Figure S1). Despite its simpler setup, our confidence-weighted logistic combination model was able to provide team complementarity not only for human-machine teams but also for machine-machine teams (Figure S1). Furthermore, our approach compared favorably to the Bayesian combination model ($t_{Welch}(20)=2.91$, $P<0.01$).
Figure 1 shows the performance of our confidence-weighted logistic combination model in the case of high noise, where most machines outperform humans. Of primary interest was whether teams including humans performed better than machine-only teams. We assessed machine-only teams comprising either one or two machines. Adding a human teammate to machine-only teams always improved the team’s performance ($t_{Welch}(22)=4.70$, $P<0.0001$).
Figure 2: Assessing Humans and LLMs using BrainBench [14]. (a) The benchmark consisted of test cases constructed from the Journal of Neuroscience abstracts. Abstracts consist of background, methods, and results. The test-taker chose which of two versions of the abstract was the original version. The altered version maintained coherency while significantly altering the results. The 100 test cases considered here were constructed by GPT-4 with human oversight and quality control. (b) An example test case. Humans were instructed to select which version of the abstract was the original by clicking on either blue or green text to select that set of options. Test cases varied in the numbers of alternatives, but a single click will choose all options of the same color. After their choice, humans indicated their confidence. LLMs chose the version of the abstract that had the lower perplexity and their confidence was assessed by the absolute difference in perplexity of the two options.
Figure 3: Conditions for effective collaboration between human experts and LLMs were satisfied. (a) When LLMs were confident in their BrainBench judgments, they were more likely to be correct. Confidence ratings were sorted into equal bins, and the mean accuracy for each bin was plotted. The positive slope of the black regression lines for Llama2 chat models (7B, 13B, and 70B) indicates well-calibrated confidence, meaning higher confidence correlated with higher accuracy. (b) LLMs surpassed human experts on BrainBench overall. Error bars represent standard error of the mean using a binomial model. (c) Item difficulty Spearman correlations among LLMs and human experts. For LLMs, difference in perplexity between incorrect and correct abstracts was used to determine the relative difficulty of test cases. Mean accuracy was used for human experts. LLMs align more with each other than humans, which implies human-machine teams will be diverse.
Bayesian integration is marked by combining judgments based on confidence ratings from team members, whether human or machine. Our confidence model can use this information as well. However, in this dataset, is confidence weighting crucial or could improved team performance simply reflect an averaging or wisdom-of-the-crowd effect? This question has not been asked with this dataset, but is straightforward to evaluate using variations of our regression approach. We found that removing confidence in our approach did not negatively impact team performance in this noisy object recognition task (Figure S2). For human-machine teams, there is no difference between our standard model and a variant that does weight by confidence ($t_{Welch}(26)=0.29$, $P=0.39$). Surprisingly, for machine-machine teams, the no-confidence model outperformed the standard ($t_{Welch}(17)=2.58$, $P<0.01$). On the one hand, these results demonstrate the versatility of our approach in that different variations of our model can assess the basis for the success of human-machine teams. On the other hand, the fact that the object recognition dataset did not require confidence-weighted integration motivates considering another dataset for which we know that confidence of humans and machines are calibrated in that higher confidence is associated with higher accuracy. The next dataset considered provides a valuable test of our standard confidence-weighted model.
Figure 4: Performance of all possible teams using the confidence-weighted logistic combination model. Adding a human to a team with one or more machines (blue points) always has a benefit (green points). Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
Calibration of confidence and classification diversity afford reliable human-LLM teaming
Our contribution also relies on previous efforts that developed BrainBench [14] to assess the capacity of humans and LLMs to predict the outcomes of neuroscience studies. The benchmark includes test cases based on abstracts from the Journal of Neuroscience. Each test case contains an original abstract and an altered version (Figure 2). The BrainBench task is to identify the correct study outcome by choosing between the original abstract and its altered counterpart.
We evaluated the conditions for effective collaboration (i.e., complementarity), namely calibration of confidence and classification diversity among team members. Both humans and LLMs were calibrated in that accuracy was positively correlated with confidence (Figure 3a). Diversity held in that LLMs and humans differed on which test items led to errors (Figure 3c). In terms of accuracy (Figure 3b), LLMs numerically surpassed humans ($t(2)=5.20$, $P<0.05$). Thus, we can consider whether humans can benefit teams consisting of machines that perform comparably or better.
Similarly as in the image classification task, we investigated whether teams including humans performed better than LLM-only teams in the prediction of neuroscience results. All 15 possible team combinations, ranging from individual teammates to a 4-way human-LLM team, were considered (Figure 4). Adding a human teammate to LLM-only teams always improved the team’s performance ($t_{Welch}(8)=8.24$, $P<0.0001$).
In the object recognition task, team performance is not affected by teammates’ confidence. Does the fluctuating confidence on a trial by trial basis matter for LLMs? Or could complementarity be achieved by just forming a weighted average of responses as in the object recognition task? We evaluated the impact of confidence by setting the magnitude of the confidence scores to 1. This setting mimics the concept of “the wisdom of the crowds,” where responses are considered without factoring in confidence levels. We find that confidence is particularly important in this knowledge-intensive task, as the overall performance diminishes when confidence scores are neglected (Figure 5). With this variant, human-LLM teams do not always surpass LLM-only teams ($t_{Welch}(7)=0.38$, $P=0.36$). The model including confidence (Figure 4) outperforms the variant without confidence (Figure 5) by a significant margin ($t_{Welch}(17)=3.16$, $P<0.01$).
We then assessed whether our regression approach would benefit from a more elaborated formulation. Eq. (3) presents a function that modulates the magnitude of a team member’s confidence rating to adjust their calibration. In an optimal setting (Figure S3), the resulting model was indistinguishable from the standard regression model ($t_{Welch}(19)=0.05$, $P=0.48$). Adding interaction terms to the confidence-weighted features does not improve the performance of either human-LLM or LLM-only teams (Figure S4). More complex model variants may perform better in other tasks, especially when there is more and less noisy data. Simpler model variants seem to be more robust to noise in our cross-validated experiments. Thus, the simplicity of our approach provides effective team collaboration.
Figure 5: Removing confidence from the logistic combination model diminishes team performance. Accuracy results on the neuroscience forecasting task with the confidence-weighted logistic combination model, where confidence scores were set to 1, i.e., $f(x)=1$ in Eq. (3). Adding a human to a team with one or more machines (blue points) does not necessarily improve performance (green points). Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model.
Our confidence-weighted logistic regression approach follows from the principles of a Bayesian combination model that fosters human-machine complementarity [13]. One questions is how well our logistic regression approach compares to the Bayesian approach beyond image classification. Our confidence-weighted regression model compared favorably (Figure 6) to the Bayesian model when evaluated on the three human-LLM teams for which the Bayesian model is intended to apply ($t_{Welch}(2)=4.08$, $P<0.05$). This success is impressive given that the regression approach takes seconds to compute on a current desktop whereas the Bayesian approach would take 12-13 days.
3 Discussion
Can humans team effectively with machines such as LLMs when the humans perform worse? We developed a confidence-weighted regression approach that can integrate judgments from any number of teammates. Using this method and testing on two forecasting benchmarks [13, 14], we found that human-machine teams achieve complementarity, that is their combined performance bests that of either teammate alone (Figures 1 and 4). Complementarity was achieved because two critical conditions were satisfied, namely confidence was well-calibrated and classification diversity held among teammates (Figure 3). Strikingly, every combination of machines benefited from adding a human to the team (Figures 1 and 4).
Figure 6: Comparison between Bayesian and confidence-weighted logistic combination models for human-LLM teams. Each data point corresponds to the average across 503 test case evaluations. Llama2 chat 7B, 13B, and 70B models were considered. Error bars represent standard error of the mean using a binomial model. The confidence-weighted logistic combination model more effectively integrated human and machine judgments.
Confidence in individual responses significantly impacted team performance in the knowledge-intensive task. The overall performance diminished when confidence scores were neglected (Figure 5). However, in the noisy object recognition task, this effect was not present (Figure S2). In both scenarios, the performance of the confidence-weighted integration of judgments from teammates was equal to or better than the weighted average of responses, i.e., “the wisdom of the crowds.” For the confidence-weighted model variants to thrive, confidence judgments need to be calibrated such that higher confidence is assocaited with higher accuracy.
Our approach was informed by a Bayesian method for combining judgments of humans and machines [13]. Our approach has a number of advantages, including ease of implementation, very fast runtime, an interpretable solution, and readily extendable to any number of teammates. Surprisingly, our confidence-weighted regression approach performed better than the Bayesian approach (Figure 6). One possibility is that the discretization of continuous confidence measures, which the Bayesian model requires, limited its performance. Perhaps an alternative formulation would perform better. Unfortunately, reformulating the Bayesian model and properly implementing it requires substantial effort and expertise. In contrast, because our confidence-weighted integration model is formulated within a regression framework, it is straightforward to extend the formulation. For example, by including a function that modulates the calibration of a teammate (Figure S3), or nonlinear relationships (e.g., polynomial terms) between confidence-weighted predictions and outcomes (Figure S4).
While we selected three LLMs with superhuman performance on BrainBench, these LLMs are not the highest performing models on this benchmark [14]. Our choice was deliberate because a vastly superior teammate may hinder complementarity. In the limit, a teammate who is never wrong does not need to be part of a team. This limiting condition may become more prevalent should LLMs continue to improve and, therefore, diminish the benefits of human-LLM teaming. For the foreseeable future, we suspect there will be tasks for which humans and LLMs can effectively team. Moreover, our method for integrating the judgments of teammates is not limited to human-LLM teams. Instead, the method is general and applies to any set of agents (natural or artificial) that can report how confident they are in their decisions.
This study explored the possibility of a collaborative approach between humans and machines for superior decision-making in classifying noisy natural images and forecasting neuroscience outcomes. Our confidence-weighted regression method effectively combined human and machine judgments because teammates fulfilled the conditions of well-calibrated confidence and classification diversity. Our results suggest there is a place for humans in teams with machines even when the machines perform better. We hope our work facilitates successful collaborations between humans and machines in addressing important challenges.
4 Methods
Datasets
Noisy LSVRC-2012 ImageNet [13]
A subset of the 2012 Large Scale Visual Recognition Challenge (LSVRC) ImageNet training set[17] was utilized. Namely, the dataset comprised 1 200 test cases (i.e., images), divided equally into 16 classes (chair, oven, knife, bottle, keyboard, clock, boat, bicycle, airplane, truck, car, elephant, bear, dog, cat, and bird). Two levels of phase noise were independently applied to distort the image dataset, low ($\Omega=80$) and high ($\Omega=125$).
BrainBench [14]
The benchmark includes test cases created either by expert neuroscientists or by prompting GPT-4 (Azure OpenAI API; version 2023-05-15) to create test cases. Since LLMs outperform humans by a large margin in both scenarios [14] and substantial differences in performance may preclude complementarity [13], we used the GPT-4 generated test cases because the performance difference between humans and LLMs was smaller for these test cases, though LLMs were still clearly superior. We considered a dataset comprising 100 machine-generated test cases. These test cases were created from abstracts in the Journal of Neuroscience published in 2023. These abstracts are categorized into five sections: Behavioral/Cognitive, Systems/Circuits, Development/Plasticity/Repair, Neurobiology of Disease, and Cellular/Molecular. Each test case contains a published abstract and an altered version produced by GPT-4. These modifications, though minimal, significantly change the results—for instance, by changing the roles of brain regions or reversing a result’s direction (e.g., from “decreases” to “increases”). The altered abstracts remain logically coherent despite the changes. The BrainBench task is to identify the correct study outcome by choosing between the original abstract and its altered counterpart.
Procedure for human participants
Noisy LSVRC-2012 ImageNet [13]
One hundred forty-five participants classified between 34 and 74 noisy images (i.e., test cases) into 16 categories. For each evaluation, participants also provided a discrete confidence level (low, medium, or high). The total number of human classifications corresponded to 7 247 for low noise images, and 7 239 for high noise images.
BrainBench [14]
One hundred seventy-one neuroscience experts were recruited to complete an online study [14]. Each participant evaluated three out of the 100 test cases. Two versions of an abstract were presented, one with the actual results and one that was altered (Figure 2). Participants chose the version they believed to be the original and rated their confidence using a slider bar. After applying several exclusion criteria, the 171 participants yielded 503 observations (2 to 9 instances per test case).
Procedure for machines
Noisy LSVRC-2012 ImageNet [13]
Five different machine classifiers pretrained for ImageNet were utilized: AlexNet [18], DenseNet161 [19], GoogleNet [20], ResNet152 [21], and VGG-19 [22]. One pass through the noisy images data (epoch) was performed during stochastic gradient training. For a given image, the classifiers produced probability confidence scores for each of the 16 classes. The class label was assigned to the class with the highest probability confidence score.
BrainBench [14]
We considered LLMs from the Llama 2 chat family with 7B, 13B, and 70B weights [23]. LLMs chose the version of the abstract with the lower perplexity (PPL). Confidence was calculated as the absolute difference in PPL between the original and altered versions of the abstract (Figure 2).
Bayesian combination model
Human and machine judgments were combined by adapting a Bayesian framework for human-machine complementarity [13]. The problem setting for combining two team members, human and machine, is as follows: Let $N$ denote the number of test cases to be analyzed with $L$ possible choices. The ground truth labels of the original abstracts are $z\in\{0,\ldots,L\}^{N}$. For the human classifier, the predicted labels $y\in\{0,\ldots,L\}^{N}$ and their corresponding confidence ratings $r\in\{0,\ldots,R-1\}^{N}$ with “0: lowest possible confidence” and “$R-1$: highest possible confidence” are given. For the machine classifier, we used the probability confidence scores $\pi\in\mathbb{R}_{+}^{N\times L}$. In the case of LLMs, $\pi=\text{Softmax}(-q)$, where $q$ is the PPL, reflecting a measure of uncertainty [14].
The first step of this model is to generate correlated probability confidence scores for human and machine classifiers using a bivariate normal distribution,
$\left(\begin{matrix}\pi_{H}\\ \pi_{M}\end{matrix}\right)\sim\mathcal{N}\left(\left(\begin{matrix}\mu_{H}\\ \mu_{M}\end{matrix}\right),\left(\begin{matrix}\sigma_{H}^{2}&\sigma_{H}\sigma_{M}\rho_{HM}\\ \sigma_{H}\sigma_{M}\rho_{HM}&\sigma_{M}^{2}\end{matrix}\right)\right)\,.$ (1)
The means of the underlying distribution, $\mu_{H}$ and $\mu_{M}$, depend whether the label of test case $i$, $z_{i}$, is correct or not, i.e.,
$\mu_{H,i,j}$ $=b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,,$
$\mu_{M,i,j}$ $=b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)\,,$
with $\mathcal{Z}_{i}=\{x\,|\,x=z_{i}\}$. Note that the scalar parameters $a_{H}$, $a_{M}$, $b_{H}$, $b_{M}$, $\sigma_{H}$, $\sigma_{M}$, and $\rho_{HM}$ in Eq. (1) are learned from data. The parameter $\rho_{HM}$ from the covariance matrix learns the correlation between the human and machine classifiers. In the case of the machine classifier, $\pi_{M}$ is compared to the empirical probability confidence scores, $\pi$. Then, for the human classifier, $\pi_{H}$ is a latent variable that is used to calculate classifications
$y_{H}\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)\,,$
where $\tau$ denotes a temperature parameter, usually small, that helps convergence [13]. The predicted classifications, $y_{H}$, are compared to actual human predictions, $y$. The confidence ratings are casted via an ordered logistic model [24] that maps the continuous probability confidence scores, $\pi_{H}$, to an ordinal confidence rating, $r_{H}$. This means,
$r_{H}\sim\text{OrderedLogistic}(\pi_{H},c,\delta)\,,$
where the parameters $c\in\mathbb{R}_{+}^{R-1}$ are the breakpoints of the intervals that map $\pi_{H}$ into $r_{H}$, and $\delta$ is a scalar that controls the sharpness of the rating probability curves. Finally, $r_{H}$ is compared to the empirical human confidence ratings, $r$.
Confidence-weighted logistic combination model
We introduce a logistic regression approach that combines the judgments of any number of teammates. The logistic combination model follows the principles of the Bayesian combination model, but is formulated within an easier to implement and extend regression framework. In its most basic form, which we consider here, each teammate is captured by a single predictor in the regression model. The value of the predictor on a trial depends on the teammate’s choice and their confidence. In particular, the magnitude of the predictor is the teammate’s confidence on that trial (i.e., confidence-weighted integration) and the sign is determined by the teammate’s choice. In general, the fitted beta weight for a teammate will reflect their accuracy and calibration. In a multiclass setting, multinomial regression is utilized.
As in the Bayesian combination model, $y$, $r$, $\pi$ are given. Thus, the logistic function for the $i$-th test case is of the form:
$p_{x}=\frac{1}{1+e^{-\beta^{\top}x}}\,,$
where $p_{x}$ is the predicted probability of the arbitrarily assigned first option, and the evidence is
$\beta^{\top}x=\beta_{I}+\beta_{H}\cdot x_{H}+\beta_{M}\cdot x_{M}\,.$ (2)
The fitted weights $\beta_{I}$, $\beta_{H}$, and $\beta_{M}$ correspond to the intercept, and human and machine teammates, respectively. The term $x_{k}$ is the signed confidence for team member $k$. For human participants, $x_{H}$ is $r_{i}$ if $y_{i}$ is the selected class label and $-r_{i}$ otherwise. Thus,
$x_{H}=r_{i}(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,,$
where $\mathbf{1}_{\mathcal{A}}(y_{i})$ is an indicator function with $\mathcal{A}=\{x\,|\,x=1\}$. Hence, $\mathbf{1}_{\mathcal{A}}(x)=1$ if $x\in\mathcal{A}$, and $\mathbf{1}_{\mathcal{A}}(x)=0$ otherwise. For machine team members, it reads
$x_{M}=\pi_{i}(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,.$
In the binary case of LLMs, we consider the PPL difference as
$x_{M}=|\Delta q_{i}|(1-2\cdot\mathbf{1}_{\mathcal{A}}(y_{i}))\,.$
Consider the case where an agent is accurate, but its confidence is not calibrated. Then, a slightly more complex version of this logistic regression can be formulated by introducing a single-parameter function for each team member. This function can either pass through the magnitude of a team member’s confidence rating or squash its magnitude toward 1 for every confidence rating, e.g.,
$f(x)=1+\frac{x-1}{1+\alpha|x-1|}\,.$ (3)
That more complex model should work better if an agent is a good predictor of outcomes but its confidence is random. In that case, $\alpha$ will be high, and the magnitude of the confidence weighting will always be $f(x)=1$ in Eq. (3). Thus, the fitted weights for logistic regression would not penalize the agent as much in the combination. For a really well-calibrated agent, $\alpha$ should be close to 0. Thus, Eq. (3) becomes the identity function, $f(x)=x$.
It turns out that for the two tasks, there was no significant difference when including a more complex model optimizing $\alpha$. However, the no confidence variant (Figure 5) was considered as a foil to our approach which values confidence.
Additionally, this model can be easily expanded by including additional terms to Eq. (2). For example, a third fitted weight could be included for an interaction term $x_{H}x_{M}$. Likewise, polynomial regression could be used to include $x_{H}^{2}$ and $x_{M}^{2}$, and corresponding fitted weights.
Cross-validation procedure
Given $M$ observations for $N$ test cases (with $M>N$), we performed a leave-one-out cross-validation (LOOCV) [25, 26] that provides the best bias-variance trade-off for small datasets. Consider the evaluation of the $i$-th test case. With this procedure, we removed all instances of test case $i$ leaving the remaining $N-1$ test cases with all their instances to train the classifier teams. For testing, we utilized all the instances of test case $i$. This was repeated for all $N$ test cases, yielding $M$ predictions. Note that for individual teammates the evaluation comprised only the testing phase. When the predicted labels of any team or individual teammate in this study were randomly shuffled, the LOOCV accuracy dropped approximately to chance level (i.e., $1/L$).
Acknowledgments
This work was supported by a NeuroData Discovery Award from the Kavli Foundation to F.Y., the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), a Royal Society Wolfson Fellowship (18302), and an AI safety grant from the Foresight Institute to B.C.L.
Author contributions
Conceptualization, F.Y. and B.C.L; methodology, F.Y., X.L., and B.C.L; investigation, F.Y. and X.L.; writing-–original draft, F.Y.; writing-–review & editing, F.Y., X.L., and B.C.L; funding acquisition, F.Y. and B.C.L; resources, F.Y. and O.V.M.; supervision, B.C.L.
Declaration of interests
The authors declare no competing interests.
Supplementary information
Implementation of Bayesian combination model
Algorithm S1 Bayesian Combination Model
1:Set priors: $\alpha=1$, $a_{M}\sim\mathcal{N}(0,10)$, $b_{M}\sim\mathcal{N}(0,10)$, $\sigma_{M}\sim\text{Uniform}(0,15)$, $a_{H}\sim\mathcal{N}(0,10)$, $b_{H}=0$, $\sigma_{H}=1$, $\rho\sim\text{Uniform}(-1,1)$, $\tau=0.05$, $c\sim\text{Uniform}(0,1)$ with $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2$, and $\delta\sim\text{Uniform}(0,100)$.
2:start training:
3:     $p\sim\text{Dirichlet}(\alpha)$
4:     $z\leftarrow\text{argmax}(p)$$\triangleright$ Compare to actual data
5:     $\mu_{M,i,j}\leftarrow b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
6:     $\mu_{H,i,j}\leftarrow b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
7:     $\pi_{M}\sim\mathcal{N}\left(\mu_{M},\,\sigma_{M}\right)$triangleright Compare to actual data
8:     $\pi_{H}\sim\mathcal{N}\left(\mu_{H}+\rho\,\sigma_{H}\left(\frac{\pi_{M}-\mu_{M}}{\sigma_{M}}\right),\,\sqrt{1-\rho^{2}}\,\sigma_{H}\right)$
9:     $y\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)$triangleright Compare to actual data
10:     $r\sim\text{OrderedLogistic}(\pi_{H},c,\delta)$triangleright Compare to actual data
11:end training
12:start testing:
13:     $p\sim\text{Dirichlet}(\alpha)$
14:     $z\leftarrow\text{argmax}(p)$
15:     $\mu_{M,i,j}\leftarrow b_{M}+(a_{M}-b_{M})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
16:     $\mu_{H,i,j}\leftarrow b_{H}+(a_{H}-b_{H})\cdot\mathbf{1}_{\mathcal{Z}_{i}}(j)$
17:     $\pi_{M}\sim\mathcal{N}\left(\mu_{M},\,\sigma_{M}\right)$triangleright Compare to actual data
18:     $\pi_{H}\sim\mathcal{N}\left(\mu_{H}+\rho\,\sigma_{H}\left(\frac{\pi_{M}-\mu_{M}}{\sigma_{M}}\right),\,\sqrt{1-\rho^{2}}\,\sigma_{H}\right)$
19:     $y\sim\text{Categorical}\left(\text{Softmax}\left(\pi_{H}/\tau\right)\right)$triangleright Compare to actual data
20:     $r\sim\text{OrderedLogistic}(\pi_{H},c,\delta)$triangleright Compare to actual data
21:end testing
Algorithm S1 illustrates the inference procedure of the Bayesian combination model, which comprises two stages: estimation of parameters (training) and prediction of label probabilities (testing). For the estimation of parameters, the given data is: true labels $z$, probability confidence scores $\pi_{M}$, and human classification $y$ and confidence ratings $r$. We assumed that all human participants shared the same set of parameters ($a_{H}$ , $b_{H}$ , $\sigma_{H}$, $c$, $\delta$, and $\tau$). In the neuroscience forecasting task, human confidence ratings on the slider bar were mapped to range between 1 and 100. A wide range is computationally expensive, thus, we aggregate it into three levels: “0: low confidence”, “1: moderate confidence”, and “2: high confidence”. Then, the aggregated confidence rating used for analysis, $r\in\{0,1,2\}^{N}$, reads
[…]
The utilized cutpoints (i.e., 33 and 66) produced a good agreement between confidence and accuracy. Among the evaluations of human participants, “low” had $63.2\%$ average accuracy ($n=174$), “moderate” had $66.5\%$ ($n=185$), and “high” had $81.9\%$ ($n=144$).
To estimate the posterior over the underlying parameters, a No-U-Turn Sampler (NUTS) for Markov chain Monte Carlo (MCMC) [27] was used with $n_{w}=500$ warmup steps, $n_{c}=3$ chains, and $n_{s}=25$ samples. For the prediction of label probabilities the given data is the same as before but $z$, that is now latent. The prior distributions were set according to [13]. The correlation between team members was $\rho\sim\mathrm{Uniform}(-1,1)$. For the LLM classifier, $a\sim\mathcal{N}(0,10)$, $b\sim\mathcal{N}(0,10)$, and $\sigma\sim\mathrm{Uniform}(0,15)$. For human classifications, $a$ was the same but $b$ and $\sigma$ were adjusted to $b=0$ and $\sigma=1$ for the purpose of identifiability [13]. In addition, we used $\delta\sim\mathrm{Uniform}(0,100)$ for the scaling parameter and uniform priors on the cutpoints, $c\sim\mathrm{Uniform}(0,1)$, with the constraint that the cutpoints are ordered (i.e. $c_{i}<c_{i+1}\;\forall\,i=1,\ldots,R-2)$. Finally, we set $\tau=0.05$ for best convergence results [13].
Figure S1: Performance of the Bayesian and confidence-weighted logistic combination models in a low noise object recognition task [13]. Accuracy results on low levels of image noise ($\Omega=80$) with the Bayesian and confidence-weighted combination models, where humans outperform machines ($t(4)=-4.77$, $P<0.01$). Each data point corresponds to the average across 7 247 image evaluations. Error bars represent standard error of the mean using a binomial model. The performance between the Bayesian and confidence-weighted combination models is indistinguishable for human-machine teams ($t_{Welch}(6)=0.49$, $P=0.32$). However, the confidence-weighted combination model presented improved results for machine-machine teams ($t_{Welch}(16)=3.75$, $P<0.001$).
Figure S2: Removing confidence from the confidence-weighted logistic combination model does not negatively impact team performance in a noisy object recognition task [13]. Accuracy results on high levels of image noise ($\Omega=125$) with the confidence-weighted logistic combination model, where confidence scores were set to 1, i.e., $f(x)=1$ in Eq. (3). Each data point corresponds to the average across 7 239 image evaluations. Error bars represent standard error of the mean using a binomial model. Similarly as in the standard scenario (Figure 1), human-machine teams surpass machine-only teams ($t_{Welch}(24)=2.80$, $P<0.01$), although with a smaller margin.
Figure S3: Optimal squashing in the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14]. Accuracy results on the neuroscience forecasting task with the confidence-weighted logistic combination model, where the parameter $\alpha$ in Eq. (3) is optimized. Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Error bars represent standard error of the mean using a binomial model. Similarly as in the standard scenario (Figure 4), human-LLM teams surpass LLM-only teams ($t_{Welch}(11)=7.20$, $P<0.0001$). The performance between standard and squashing scenarios is, however, indistinguishable ($t_{Welch}(19)=0.05$, $P=0.48$).
Figure S4: Adding interaction terms to the confidence-weighted logistic combination model does not improve overall team performance in the neuroscience forecasting task [14]. Accuracy results of the confidence-weighted logistic combination model on the neuroscience forecasting task as a function of the degree of the polynomial expansion of the features. Llama2 chat 7B, 13B, and 70B models were considered. Each data point corresponds to the average across 503 test case evaluations. Only interactions were considered, as the results including pure terms is identical. Team performance is independent of interaction terms.